OpenStudy (anonymous):

a group of students decide to have custom t-shirts made to promote their fund-raiser. Prices for the t-shirts are shown below. there is also $3.5 shipping and handling charge per oder. Assume that the students have a budget of $200 for the t-shirts and that the shirts must be pirchased in packs of ten. Long-sleeve blue:$65 per 10 shirts long-sleeve white: $55 per 10 shirts short-sleeve blue: $50 per 10 shirts short-sleeve white: $45 per 10 shirts. How many long sleeve, blue t-shirts can the students buy? write and solve an inequality to find the answer.

OpenStudy (texaschic101):

p = packs of 10 Long sleeve t-shirts 3.50 + 65p >= 200 --- subtract 3.50 from both sides 65p >= 200 - 3.50 65p >= 196.50 -- divide both sides by 65 p >= 196.50/65 p >= 3.02 rounds to 3 p >= 3 packs (each pack containing 10 shirts) or p >= 30 long sleeve blue t-shirts (because each pack contains 10 shirts...so that makes 30 shirts)